OpenStudy (anonymous):

simplify the radical expression \[sqrt{252}\]

OpenStudy (anonymous):

\[\sqrt{252} = \sqrt{2 \times 2 \times 3 \times 3 \times 7}\] Take out the number you can take out of the square root brackets..

OpenStudy (anonymous):

Thanks!

OpenStudy (radar):

\[\sqrt{252}=\sqrt{36\times7}=\sqrt{36}\sqrt{7}=\]You can take it from there.